OpenStudy (anonymous):

The thickness of a kerosene spill is 1.2mm. The spill is .79 miles in diameter. Kerosene has a density of .82g/cc. Assuming that the spill is uniform in thickness and is shaped in a circle, how much kerosene, volume and mass, is spilled?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

I have to find the mass in tons and the volume in gallons. I have a few numbers but I messed up on the conversions

OpenStudy (anonymous):

For example, to find the volume of the shape I used the formula of \[v = \pi r^2 h\]

OpenStudy (anonymous):

And did half of .79 do get the r (radius)

OpenStudy (phi):

I think I would work in metric to get the volume and mass then convert

OpenStudy (anonymous):

Ok, is the process I did to find the volume correct though? Or should I use another formula?

OpenStudy (anonymous):

Because I know that once I have the volume I can use that and the density to find the mass, and then carry on to my conversions

OpenStudy (phi):

yes, that is the correct formula the hard part is keeping the units consistent. I would work with cm and g , and write the numbers in scientific notation

OpenStudy (anonymous):

Yeah, that's what I did. I converted the 1.2 mm into .12 cm

OpenStudy (anonymous):

But how would I do the .79 miles? Would I convert the diameter into cm or the radius?

OpenStudy (phi):

google 1 mile in meters you get 1609.344 meters in cm that is 160934.4 cm per mile that is one mile. the radius is 0.5*0.79 miles if we write cm per miles as a fraction cm/mile we want the miles to cancel, so the right thing to do is \[ 0.5 \cdot 0.79\ \cancel{mile} \cdot \frac{160934.4 \ cm}{1 \cancel{mile} } \]

OpenStudy (anonymous):

Um ok... where did you get the .5 from?

OpenStudy (anonymous):

Oh never mind, that's just 1/2 right?

OpenStudy (phi):

½ of the diameter

OpenStudy (anonymous):

Yeah.. Okay let me work on it really fast

OpenStudy (anonymous):

Ok, I got 63569.088

OpenStudy (phi):

or about 6.3569e4 cm

OpenStudy (anonymous):

I just did .5 x .79 = .395 and then .395 x 160934.4 = 6.35

OpenStudy (anonymous):

What is the little e?

OpenStudy (phi):

that is how I write 6.3569 * 10^4 the little "e" is a reminder that the next number is the exponent of (an understood) 10

OpenStudy (phi):

I assume you know that 63569 can be written as 6.3569 * 10^4 ? or using my notation 6.3569e4 now find the volume v= pi r^2 h = 3.14* (6.3569e4)^2 * 0.12 cm^3

OpenStudy (phi):

you don't type the e into the calculator , It's only to make typing easier for me.

OpenStudy (phi):

or type into google 3.14* (63569)^2 * 0.12 =

OpenStudy (phi):

the volume in cm^3 is 1.2492 * 10^9 cm^3 now multiply by 0.82 g/cm^3 \[ 1.2492 \cdot 10^9 \ \cancel{ cm^3} \cdot0.82 \frac{g}{ \cancel{ cm^3}} \]

OpenStudy (anonymous):

Oh Ok! Sorry I was eating dinner

OpenStudy (anonymous):

And then do I have to mke any other conversions?

OpenStudy (anonymous):

Once I find the mass? I know it has to be in tons and the volume in gallons